# LVAgent: Long Video Understanding by Multi-Round Dynamical Collaboration of MLLM Agents

### v4

@@ 1 Introduction
Long videos are crucial in Internet data and important in fields like healthcare, education, and entertainment [37, 14, 35]. In the current data-driven era, correctly and efficiently understanding long videos holds great significance. Despite being pretrained on large multi-modal datasets, Multi-modal Large Language Models (MLLMs) [38, 45, 47, 30, 56, 57, 42, 60, 43] struggle with modeling the long-term temporal context of long videos (from minutes to hours long). Feeding numerous video frames (e.g., at 1 frame/s) directly into a single MLLM incurs high computational costs and much redundant information, yielding poor performance. Taking Figure 3 as an example, we input the whole video into Gemini 1.5Pro [44], GPT-4o [38], Qwen2-VL [47] and InternVL-2.5 [4], and all of these models fail to provide the correct answer due to the redundant confusing information of the video, as shown with colored text.
Agent-based approaches [8, 49, 16, 51, 58] hold great potential in long video understanding, because agents can simplify difficult problems and autonomously invoke a variety of tools to extract key information from long videos to assist in long video understanding. However, there is still significant room for improvement in current agent-based methods for long video tasks. Some methods use CLIP [39] to retrieve key frames for long video understanding [49, 58, 63, 55]. However, CLIP struggles to retrieve long-term temporal information and its pretrained data has a domain gap with long videos [46, 59]. Some approaches use external tools (e.g., memory bank [8, 13], RAG [21, 2, 16]) to aid MLLMs in long video understanding. Even with the support of tools, relying on a single MLLM to answer questions about long-form videos offers only a partial comprehension of the video content, leading to limited performance. In conclusion, there are two main challenges in long video understanding. The first challenge is how to better retrieve the video clip with key information according to the query. The second challenge is how to use multi-agent collaboration to better understand long videos.
@@
To address these challenges, we put forward a four-step dynamic collaborative paradigm for agents, namely Selection, Perception, Action, and Reflection. (1) Selection. We pre-select the optimal agent team from an agent library housing mainstream MLLMs like Qwen2-VL [47] and InternVL-2.5 [4]. With selection process, only the most suitable models are involved in the subsequent processes, thus saving computational resources and enhancing efficiency. (2) Perception. Regarding the challenge of CLIP in retrieving long video temporal info, we propose an adaptive frame extraction method and design an effective retrieval approach for long videos. This pipeline is designed to enhance the coverage of crucial temporal chunks without sacrificing computational efficiency. (3) Action. After retrieving the query-related video clips, agents answer long video questions and provide the reasons. (4) Reflection. In order to enable better collaboration among multiple agents for a better understanding of long videos, we introduce a reflection process to evaluate each agent’s performance in every discussion round. Based on performance, agents are scored and filtered, enabling dynamic collaboration for each query. This reflection process helps facilitate better collaboration, rule out wrong inferences, and reach a consensus. Throughout these four processes, agents refine answers by discussing, retrieving key video chunks, and integrating information with other pre-selected agents. We validate the effectiveness on four long video understanding datasets: EgoSchema [34], VideoMME [12], MLVU [64], LongVideoBench [53]. Experiments show that LVAgent achieves an accuracy of over 80% on mainstream long video tasks. Our contributions can be summarized as threefold:
@@
  • ␣Second, LVAgent is a novel multi-round and multi-step dynamical collaboration pipeline of MLLM Agents. LVAgent enables multi MLLM agents to process and integrate long-term temporal context dynamically for more comprehensive and robust long video understanding.
@@ 2 Related Work
Recent advancements in multimodal large language models (MLLMs), such as [47, 24, 62, 25, 20, 57, 40], have significantly advanced video understanding, particularly in temporal reasoning. However, processing long videos remains a critical challenge due to the inherent redundancies in temporal sequences, which result in an overwhelming number of visual tokens. To address this, current approaches primarily focus on two strategies: (1) extending the context length of LLMs [27, 45, 48, 56] and (2) compressing video tokens [43, 60, 52, 26, 11, 42, 57]. For context extension, methods like LongVA [61] fine-tune language models to expand their context windows, successfully generalizing short-video understanding to longer videos. However, context extension often overlooks redundant information and computational overhead caused by the increase in the number of frames during inference. On the other hand, token compression techniques, such as MovieChat [43], reduce redundancy by merging similar visual tokens. While this improves efficiency, it often sacrifices fine-grained details, leading to suboptimal performance in complex scenarios.
@@
Agent-based methods [23, 7, 50, 17, 41, 54, 1] have the ability to decompose complex tasks. Currently, there are mainly three paradigms for agents in solving long video tasks. First, some agent methods employ retrieval methods such as CLIP [39] to identify the video regions most relevant to the question [63, 55, 49]. However, CLIP-based approach suffers from a domain gap when dealing with long videos. Second, the Retrieval Augmented Generation (RAG) method [16] leverages external tools for information extraction, such as memory banks [8, 13] and search engines [21, 2], for a more comprehensive understanding. However, the accuracy of the RAG method is also limited. Third, the Chain-of-thought (CoT) method is adopted to enhance the model’s reasoning ability to answer questions correctly [51, 10]. While agents show promise for long video tasks, current agent-based methods rely only on a single MLLM for such questions, lacking a dynamic framework for multiple MLLMs to answer, discuss, and collaborate. This ineffective collaboration results in an incomplete and partial understanding of long videos. There is still much room for improvement in accuracy. Specifically, none of the existing agent paradigms has outperformed the best Multimodal Large Language Models (MLLMs) [47, 38, 62, 44] on the long video understanding tasks [34, 12, 65, 53]. Fortunately, multi-round dynamical collaboration pipeline [29] demonstrated markable performance in NLP tasks. Inspired by this, we introduce our multi-round dynamical collaboration of MLLM agents for long video understanding.
@@ 3.1 Selection
To answer the question about long videos, we leverage the popular MLLMs to establish an Agent Library, which encompasses currently popular MLLMs, i.e. Qwen2-VL [47], InternVL-2.5 [4], LongVU [40], and LLaVA-Video (LV) [62]. In multi-agent systems, agents often vary in performance across tasks and domains, with underperforming agents disrupting collaboration and increasing computational costs. To address this, we propose an agent pre-selection method based on pseudo-label voting to efficiently identify the optimal agent team for a task as shown in Figure 4.
@@ 3.2 Perception
In stage 3, we perform detail video chunks retrieval. Through ablation experiments, we divide the video into six equal chunks $\{chk_{1},chk_{2},...,chk_{6}\}$. For each chunk, we randomly sample 16 frames to form a frame set $S_{frame}={f_{1},f_{2},...,f_{6}}$. Then, we use ASP-CLIP model[3] finetuned on our collected LongVR dataset to calculate the CLIP score for each video chunk.
@@
To enhance the retrieval ability of ASP-CLIP to better assist with long video tasks, we compile a diverse dataset for fine-tuning the ASP-CLIP model, consisting of 82k video clips from five datasets, as shown in Table 8 in the appendix. This collection covers a wide range of scenes, themes, and video quality, providing a robust foundation for training a video retrieval model. We extract long videos from the datasets of ActivityNet [19], OpenVid-1M [36], ViTT [15], MovieChat-Caption [43], and Youcook2 [66] for the retrieval task, as depicted in Figure 6. First, we crop the videos clips with captions. If a video has been pre-annotated into multiple clips, we take these clips along with their captions as components of our dataset. Second, to better suit the long video scenario, we filter out video clips shorter than 5 seconds or longer than 12 minutes. Only clips of appropriate length and their captions are retained. Finally, considering that the context length of CLIP is limited, to prevent information truncation in the clip-text encoder due to overly long captions, we use Qwen2-VL [47] for caption optimization. Additionally, if the number of tokens in a caption is less than 20, we also filter out this data, as such short captions are insufficient to summarize the content of a relatively long video. After manually screening the captions, the resulting long video dataset has an average video duration of 145.6 seconds. The captions in the dataset have an average length of 71 tokens. In total, we assemble a long video retrieval dataset with 82K samples. The details of dataset constructions and data recipes can be found in the supplementary document. We finetune ASP-CLIP [3] on LongVR and achieved significant improvements.
@@
GPT-4o [38] 72.2 66.7 64.6 80.0 / 82.8 70.3 / 76.6 65.3 / 72.1 71.9 / 77.2
Gemini 1.5Pro [44] 71.1 64.0 - 81.7 / 84.5 74.3 / 81.0 67.4 / 77.4 75.0 / 81.3
ShareGPT4Video-8B [2] - 39.7 46.4 48.3 / 53.6 36.3 / 39.3 35.0 / 37.9 39.9 / 43.6
VideoChat2-7B [25] 56.7 39.3 47.9 48.3 / 52.8 37.0 / 39.4 33.2 / 39.2 39.5 / 43.8
InternVL-2.5-8B [4] - 60.0 68.9 - - - 64.2 / 66.9
Qwen2-VL-7B [47] 66.7 55.6 - - - - 63.3 / 69.0
LongVA-7B [61] - 51.3 58.8 61.1 / 61.6 50.4 / 53.6 46.2 / 47.6 52.6 / 54.3
LongVU-7B [40] 67.6 - 65.4 - - - / 59.5 - / 60.6
LLaVA-Video-7B [62] 57.3 58.2 70.8 - - - 63.3 / 69.7
Aria-28B [22] - 64.2 72.3 67.9 / 78.3 67.0 / 71.7 58.8 / 66.3 67.6 / 72.1
Oryx-34B [28] - - 70.6 77.3 / 80.6 65.3 / 74.3 59.3 / 69.9 67.3 / 74.9
Qwen2-VL-72B [47] 77.9 - - 80.1 / 82.2 71.3 / 76.8 62.2 / 74.3 71.2 / 77.8
LLAVA-Video-72B [62] 65.6 64.9 - 81.4 / 82.8 68.9 / 75.6 61.5 / 72.5 70.6 / 76.9
InternVL-2.5-78B [4] - 63.6 75.7 82.8 / 83.2 70.9 / 74.1 62.6 / 64.8 72.1 / 74.0
LLaVA-OneVision-72B [20] 62.0 63.2 68.0 76.7 / 79.3 62.2 / 66.9 60.0 / 62.4 66.3 / 69.6
VideoLLaMA-2-72B [6] 63.9 - 45.6 69.8 / 72.0 59.9 / 63.0 57.6 / 59.0 62.4 / 64.7
VideoAgent [49] 54.1 - - - - - -
VideoTree [51] 61.1 - - - - - -
VCA [58] - 41.3 - - - - -
DrVideo [33] 61.0 - - - - - 51.7/71.7
VideoRAG-72B [32] - 65.4 73.8 81.1 / - 72.9 / - 73.1 / - 75.7 / -
@@
Table 1: Comparison on Mainstream Long Video Understanding Tasks. The Video-MME results are presented in the format “w/o subs / w/ subs". We are the first agent-based method to achieve over 80% accuracy on each long video understanding dataset
@@ 4.1 Benchmark
We evaluate LVAgent on four long-term video question-answering benchmarks. EgoSchema [34] consists of over 5,031 human curated question answer pairs derived from Ego4D. EgoSchema requires the correct answer for each question based on a three-minute video clip. VideoMME [12] includes 900 videos with a total of 254 hours and 2,700 question-answer pairs. The video durations range from 11 seconds to 1 hour, with an average duration of 1024 seconds. MLVU [64] includes 1,730 videos and 2,174 questions in 9 categories, specifically designed for understanding long videos. The videos last between 3 minutes and over 2 hours, with an average length of 930 seconds. LongVideoBench [53] (LVBench) includes 1,337 QA pairs in val split with their subtitles in various themes. The videos cover four progressive duration groups: 8-15 seconds, 15-60 seconds, 3-10 minutes, and 15-60 minutes, and the average duration is 473 seconds.
@@
Qwen2-VL[47] 568 90.5s 55.6 71.2 / 77.8
GPT-4o[38] 384 153.6s 66.7 71.9 / 77.2
Gemini-1.5-Pro[44] 568 227.2s 64.0 75.0 / 81.3
@@ 4.2 Implication Details
The model libraries we construct include Qwen2-VL-7B/72B [47], LLaVA-Video-72B [62], LongVU-7B [40], InternVL-2.5-7B/78B [4], Oryx [28] and Aria [22]. To avoid randomness in the results, when processing, we adjust the temperature coefficient of all MLLMs to 0.01 and uniformly set the maximum number of new tokens to be generated to 168, thus preventing the generation of excessive redundant information. For the retrieval model, we use ASP-CLIP [3] model for better temporal context modeling. We initialize the visual and textual encoders with the pretrained CLIP, following the previous methods [31, 9, 5]. To construct temporal information, we utilize a four-layer temporal transformer featuring 8 attention heads. The initial learning rate is configured as 1e-7 for both the text encoder and video encoder, while for other parameters, it is set to 1e-4. Our ASP-CLIP is optimized using Adam [18], with a batch size of 64. The model is trained for 10 epochs with a cosine learning rate schedule. The training process and our experiments on LVAgent are based on Pytorch with 8 A800-80G GPUs.

### v5

@@ 1 Introduction
Long videos are crucial in Internet data and important in fields like healthcare, education, and entertainment [videoinedu, videoinmed, videoinent]. In the current data-driven era, correctly and efficiently understanding long videos holds great significance. Despite being pretrained on large multi-modal datasets, Multi-modal Large Language Models (MLLMs) [gpt4o, gemini, qwen2vl, liu2024nvila, longvila, timesuite, videoxl, videollama, moviechat, chen2022low, chen2025g] struggle with modeling the long-term temporal context of long videos (from minutes to hours long). Feeding numerous video frames (e.g., at 1 frame/s) directly into a single MLLM incurs high computational costs and much redundant information, yielding poor performance. Taking Figure 3 as an example, we input the whole video into Gemini 1.5Pro [gemini15], GPT-4o [gpt4o], Qwen2-VL [qwen2vl] and InternVL-2.5 [internvl2.5], and all of these models fail to provide the correct answer due to the redundant confusing information of the video, as shown with colored text.
Agent-based approaches [videoagentmemory, videoagentrepeat, videorag, videotree, vca, chen2025vragent, chen2025videochat, yue2025uniflow, chen2025top] hold great potential in long video understanding, because agents can simplify difficult problems and autonomously invoke a variety of tools to extract key information from long videos to assist in long video understanding. However, there is still significant room for improvement in current agent-based methods for long video tasks. Some methods use CLIP [clip] to retrieve key frames for long video understanding [videoagentrepeat, vca, zhao2024longagent, xie2024openagents]. However, CLIP struggles to retrieve long-term temporal information and its pretrained data has a domain gap with long videos [videoclipxl, longclip]. Some approaches use external tools (e.g., memory bank [videoagentmemory, mallm], RAG [searchlvlm, chen2024sharegpt4video, videorag]) to aid MLLMs in long video understanding. Even with the support of tools, relying on a single MLLM to answer questions about long-form videos offers only a partial comprehension of the video content, leading to limited performance. In conclusion, there are two main challenges in long video understanding. The first challenge is how to better retrieve the video clip with key information according to the query. The second challenge is how to use multi-agent collaboration to better understand long videos.
@@
To address these challenges, we put forward a four-step dynamic collaborative paradigm for agents, namely Selection, Perception, Action, and Reflection. (1) Selection. We pre-select the optimal agent team from an agent library housing mainstream MLLMs like Qwen2-VL [qwen2vl] and InternVL-2.5 [internvl2.5]. With selection process, only the most suitable models are involved in the subsequent processes, thus saving computational resources and enhancing efficiency. (2) Perception. Regarding the challenge of CLIP in retrieving long video temporal info, we propose an adaptive frame extraction method and design an effective retrieval approach for long videos. This pipeline is designed to enhance the coverage of crucial temporal chunks without sacrificing computational efficiency. (3) Action. After retrieving the query-related video clips, agents answer long video questions and provide the reasons. (4) Reflection. In order to enable better collaboration among multiple agents for a better understanding of long videos, we introduce a reflection process to evaluate each agent’s performance in every discussion round. Based on performance, agents are scored and filtered, enabling dynamic collaboration for each query. This reflection process helps facilitate better collaboration, rule out wrong inferences, and reach a consensus. Throughout these four processes, agents refine answers by discussing, retrieving key video chunks, and integrating information with other pre-selected agents. We validate the effectiveness on four long video understanding datasets: EgoSchema [egoschema], VideoMME [videomme], MLVU [MLVU], LongVideoBench [wu2024longvideobench]. Experiments show that LVAgent achieves an accuracy of over 80% on mainstream long video tasks. Our contributions can be summarized as threefold:
@@
  •  ̵Second, LVAgent is a novel multi-round and multi-step dynamical collaboration pipeline of MLLM Agents. LVAgent enables multi MLLM agents to process and integrate long-term temporal context dynamically for more comprehensive and robust long video understanding.
@@ 2 Related Work
Recent advancements in multimodal large language models (MLLMs), such as [qwen2vl, videochat, llavavideo, mvbench, llavaonevision, timesuite, longvu], have significantly advanced video understanding, particularly in temporal reasoning. However, processing long videos remains a critical challenge due to the inherent redundancies in temporal sequences, which result in an overwhelming number of visual tokens. To address this, current approaches primarily focus on two strategies: (1) extending the context length of LLMs [lwm, gemini, longllava, longvila] and (2) compressing video tokens [moviechat, videollama, longvlm, llamavid, videoccam, videoxl, timesuite]. For context extension, methods like LongVA [longva] fine-tune language models to expand their context windows, successfully generalizing short-video understanding to longer videos. However, context extension often overlooks redundant information and computational overhead caused by the increase in the number of frames during inference. On the other hand, token compression techniques, such as MovieChat [moviechat], reduce redundancy by merging similar visual tokens. While this improves efficiency, it often sacrifices fine-grained details, leading to suboptimal performance in complex scenarios.
@@
Agent-based methods [li2023camel, debate1-mit, multi-persona, blender, reflexion, autogen, pca] have the ability to decompose complex tasks. Currently, there are mainly three paradigms for agents in solving long video tasks. First, some agent methods employ retrieval methods such as CLIP [clip] to identify the video regions most relevant to the question [zhao2024longagent, xie2024openagents, videoagentrepeat, chen2025super]. However, CLIP-based approach suffers from a domain gap when dealing with long videos. Second, the Retrieval Augmented Generation (RAG) method [videorag] leverages external tools for information extraction, such as memory banks [videoagentmemory, mallm] and search engines [searchlvlm, chen2024sharegpt4video], for a more comprehensive understanding. However, the accuracy of the RAG method is also limited. Third, the Chain-of-thought (CoT) method is adopted to enhance the model’s reasoning ability to answer questions correctly [videotree, videoofthought]. While agents show promise for long video tasks, current agent-based methods rely only on a single MLLM for such questions, lacking a dynamic framework for multiple MLLMs to answer, discuss, and collaborate. This ineffective collaboration results in an incomplete and partial understanding of long videos. There is still much room for improvement in accuracy. Specifically, none of the existing agent paradigms has outperformed the best Multimodal Large Language Models (MLLMs) [qwen2vl, gpt4o, llavavideo, gemini15] on the long video understanding tasks [egoschema, videomme, zhou2024mlvu, wu2024longvideobench]. Fortunately, multi-round dynamical collaboration pipeline [liu2024dynamic] demonstrated markable performance in NLP tasks. Inspired by this, we introduce our multi-round dynamical collaboration of MLLM agents for long video understanding.
@@ 3.1 Selection
To answer the question about long videos, we leverage the popular MLLMs to establish an Agent Library, which encompasses currently popular MLLMs, i.e. Qwen2-VL [qwen2vl], InternVL-2.5 [internvl2.5], LongVU [longvu], and LLaVA-Video (LV) [llavavideo]. In multi-agent systems, agents often vary in performance across tasks and domains, with underperforming agents disrupting collaboration and increasing computational costs. To address this, we propose an agent pre-selection method based on pseudo-label voting to efficiently identify the optimal agent team for a task as shown in Figure 4.
@@ 3.2 Perception
In stage 3, we perform detail video chunks retrieval. Through ablation experiments, we divide the video into six equal chunks $\{chk_{1},chk_{2},...,chk_{6}\}$. For each chunk, we randomly sample 16 frames to form a frame set $S_{frame}={f_{1},f_{2},...,f_{6}}$. Then, we use ASP-CLIP model[asp] finetuned on our collected LongVR dataset to calculate the CLIP score for each video chunk.
@@
To enhance the retrieval ability of ASP-CLIP to better assist with long video tasks, we compile a diverse dataset for fine-tuning the ASP-CLIP model, consisting of 82k video clips from five datasets, as shown in Table 8 in the appendix. This collection covers a wide range of scenes, themes, and video quality, providing a robust foundation for training a video retrieval model. We extract long videos from the datasets of ActivityNet [activitynet], OpenVid-1M [openvid], ViTT [vitt], MovieChat-Caption [moviechat], and Youcook2 [youcook2] for the retrieval task, as depicted in Figure 6. First, we crop the videos clips with captions. If a video has been pre-annotated into multiple clips, we take these clips along with their captions as components of our dataset. Second, to better suit the long video scenario, we filter out video clips shorter than 5 seconds or longer than 12 minutes. Only clips of appropriate length and their captions are retained. Finally, considering that the context length of CLIP is limited, to prevent information truncation in the clip-text encoder due to overly long captions, we use Qwen2-VL [qwen2vl] for caption optimization. Additionally, if the number of tokens in a caption is less than 20, we also filter out this data, as such short captions are insufficient to summarize the content of a relatively long video. After manually screening the captions, the resulting long video dataset has an average video duration of 145.6 seconds. The captions in the dataset have an average length of 71 tokens. In total, we assemble a long video retrieval dataset with 82K samples. The details of dataset constructions and data recipes can be found in the supplementary document. We finetune ASP-CLIP [asp] on LongVR and achieved significant improvements.
@@
GPT-4o [gpt4o] 72.2 66.7 64.6 80.0 / 82.8 70.3 / 76.6 65.3 / 72.1 71.9 / 77.2
Gemini 1.5Pro [gemini15] 71.1 64.0 - 81.7 / 84.5 74.3 / 81.0 67.4 / 77.4 75.0 / 81.3
ShareGPT4Video-8B [chen2024sharegpt4video] - 39.7 46.4 48.3 / 53.6 36.3 / 39.3 35.0 / 37.9 39.9 / 43.6
VideoChat2-7B [mvbench] 56.7 39.3 47.9 48.3 / 52.8 37.0 / 39.4 33.2 / 39.2 39.5 / 43.8
InternVL-2.5-8B [internvl2.5] - 60.0 68.9 - - - 64.2 / 66.9
Qwen2-VL-7B [qwen2vl] 66.7 55.6 - - - - 63.3 / 69.0
LongVA-7B [longva] - 51.3 58.8 61.1 / 61.6 50.4 / 53.6 46.2 / 47.6 52.6 / 54.3
LongVU-7B [longvu] 67.6 - 65.4 - - - / 59.5 - / 60.6
LLaVA-Video-7B [llavavideo] 57.3 58.2 70.8 - - - 63.3 / 69.7
Aria-28B [aria] - 64.2 72.3 67.9 / 78.3 67.0 / 71.7 58.8 / 66.3 67.6 / 72.1
Oryx-34B [liu2024oryx] - - 70.6 77.3 / 80.6 65.3 / 74.3 59.3 / 69.9 67.3 / 74.9
Qwen2-VL-72B [qwen2vl] 77.9 - - 80.1 / 82.2 71.3 / 76.8 62.2 / 74.3 71.2 / 77.8
LLAVA-Video-72B [llavavideo] 65.6 64.9 - 81.4 / 82.8 68.9 / 75.6 61.5 / 72.5 70.6 / 76.9
InternVL-2.5-78B [internvl2.5] - 63.6 75.7 82.8 / 83.2 70.9 / 74.1 62.6 / 64.8 72.1 / 74.0
LLaVA-OneVision-72B [llavaonevision] 62.0 63.2 68.0 76.7 / 79.3 62.2 / 66.9 60.0 / 62.4 66.3 / 69.6
VideoLLaMA-2-72B [videollama2] 63.9 - 45.6 69.8 / 72.0 59.9 / 63.0 57.6 / 59.0 62.4 / 64.7
VideoAgent [videoagentrepeat] 54.1 - - - - - -
VideoTree [videotree] 61.1 - - - - - -
VCA [vca] - 41.3 - - - - -
DrVideo [drvideo] 61.0 - - - - - 51.7/71.7
VideoRAG-72B [videoragvideofeat] - 65.4 73.8 81.1 / - 72.9 / - 73.1 / - 75.7 / -
@@
Table 1: Comparison on Mainstream Long Video Understanding Tasks. The Video-MME results are presented in the format “w/o subs / w/ subs”. We are the first agent-based method to achieve over 80% accuracy on each long video understanding dataset
@@ 4.1 Benchmark
We evaluate LVAgent on four long-term video question-answering benchmarks. EgoSchema [egoschema] consists of over 5,031 human curated question answer pairs derived from Ego4D. EgoSchema requires the correct answer for each question based on a three-minute video clip. VideoMME [videomme] includes 900 videos with a total of 254 hours and 2,700 question-answer pairs. The video durations range from 11 seconds to 1 hour, with an average duration of 1024 seconds. MLVU [MLVU] includes 1,730 videos and 2,174 questions in 9 categories, specifically designed for understanding long videos. The videos last between 3 minutes and over 2 hours, with an average length of 930 seconds. LongVideoBench [wu2024longvideobench] (LVBench) includes 1,337 QA pairs in val split with their subtitles in various themes. The videos cover four progressive duration groups: 8-15 seconds, 15-60 seconds, 3-10 minutes, and 15-60 minutes, and the average duration is 473 seconds.
@@
Qwen2-VL[qwen2vl] 568 90.5s 55.6 71.2 / 77.8
GPT-4o[gpt4o] 384 153.6s 66.7 71.9 / 77.2
Gemini-1.5-Pro[gemini15] 568 227.2s 64.0 75.0 / 81.3
@@ 4.2 Implication Details
The model libraries we construct include Qwen2-VL-7B/72B [qwen2vl], LLaVA-Video-72B [llavavideo], LongVU-7B [longvu], InternVL-2.5-7B/78B [internvl2.5], Oryx [liu2024oryx] and Aria [aria]. To avoid randomness in the results, when processing, we adjust the temperature coefficient of all MLLMs to 0.01 and uniformly set the maximum number of new tokens to be generated to 168, thus preventing the generation of excessive redundant information. For the retrieval model, we use ASP-CLIP [asp] model for better temporal context modeling. We initialize the visual and textual encoders with the pretrained CLIP, following the previous methods [ref5, ref6, ref7]. To construct temporal information, we utilize a four-layer temporal transformer featuring 8 attention heads. The initial learning rate is configured as 1e-7 for both the text encoder and video encoder, while for other parameters, it is set to 1e-4. Our ASP-CLIP is optimized using Adam [ref43], with a batch size of 64. The model is trained for 10 epochs with a cosine learning rate schedule. The training process and our experiments on LVAgent are based on Pytorch with 8 A800-80G GPUs.